Let f(x, y,z) = e √ z−x 2−y 2 . (a) Evaluate f(2,−1,6) (b) Find the domain of f (c) Find the range of f


1
Expert's answer
2021-06-29T11:06:13-0400
"f(x,y,z)=e^{\sqrt{z-x^2-y^2}}"

a)

"f(2,-1,6)=e^{\sqrt{6-(2)^2-(-1)^2}}=e"

b)


"z-x^2-y^2\geq 0"

Domain: "z-x^2-y^2\geq 0\ or\ z\geq x^2+y^2." The set of points that lie on the elliptic paraboloid "z=x^2+y^2" or inside it.


c)

"\sqrt{z-x^2-y^2}\geq0=>e^{\sqrt{z-x^2-y^2}}\geq 1"

Range: "[1,\infin)."